# Markovian dynamics for a quantum/classical system and quantum trajectories

### v1

@@ 3.3 A quasi-free dynamics
In [13] the structure of the most general quasi-free hybrid dynamical semigroup has been obtained. The “quasi-free” requirement means to map hybrid Weyl operators into a quantity proportional to the hybrid Weyl operators [51]. In this case, the formal generator involves unbounded operators on the Hilbert space of the quantum component; so, the semigroup found in [13] is not in the class studied here. However, the comparison of the two formal generators give the hints to construct the SDE formulation of the quasi-free dynamics.
@@
$\mathscr{U}=\Xi_{1}\oplus\Xi_{0},\qquad\Xi_{1}=\mathbb{R}^{2n},\qquad\Xi_{0}=\mathbb{R}^{s},\qquad\mathsf{U}_{0}=\mathbb{S}=\{u\in\mathscr{U}:\left|u\right|\leq 1\},\qquad a_{j}^{k}(x)=a_{j}^{k}\in\mathbb{R},$
$c_{i}(x)=\alpha^{0}_{i}+\sum_{j=1}^{s}x_{j}Z^{00}_{ji},\quad\alpha^{0}_{i}\in\mathbb{R}\quad Z^{00}_{ji}\in\mathbb{R},\quad i=1,\ldots,s,\qquad g(x,u)=g(u)=P_{0}u\quad g_{i}(u)=u_{2n+i},$
@@
$\mathrm{d}X_{i}(t)=\alpha^{0}_{i}\mathrm{d}t+\sum_{j=1}^{s}X_{j}(t_{-})Z^{00}_{ji}\mathrm{d}t+\sum_{k=1}^{d}a_{i}^{k}\mathrm{d}W_{k}(t)\\ {}+\int_{\mathbb{S}}g_{i}(u)\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)+\int_{\mathscr{U}\setminus\mathbb{S}}g_{i}(u)\Pi(\mathrm{d}u,\mathrm{d}t),\qquad i=1,\ldots,s.$ (39)
@@
$H_{\rm q}=\beta^{\mathtt{T}}R+\frac{1}{2}\,R^{\mathtt{T}}P_{1}\sigma D^{11}\sigma^{\mathtt{T}}P_{1}R,\quad\beta\in\Xi_{1},\qquad H_{x}=x^{\mathtt{T}}P_{0}ZP_{1}R=\sum_{i=1}^{s}\sum_{j=1}^{2n}x_{i}Z^{01}_{ij}R_{j}.$ (40d)
@@
$\mathcal{K}[a\otimes f](x)=f(x)\Big{\{}\mathrm{i}[H_{\rm q}+H_{x},a]+\sum_{k=1}^{d}\sum_{r,l=1}^{2n}d^{k}_{l}\overline{d_{r}^{k}}\left((\sigma R)_{r}a(\sigma R)_{l}-\frac{1}{2}\left\{(\sigma R)_{r}(\sigma R)_{l},a\right\}\right)\\ {}+a\sum_{i=1}^{s}\left(\alpha^{0}_{i}+\sum_{j=1}^{s}x_{j}Z^{00}_{ji}\right)\frac{\partial f(x)}{\partial x_{i}}+a\sum_{i,j=1}^{s}\frac{1}{2}\,\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}a_{i}^{k}a_{j}^{k}\\ {}+\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\sum_{k=1}^{d}a_{i}^{k}\sum_{l=1}^{2n}\left(\mathrm{i}(\sigma R)_{l}\overline{d_{l}^{k}}a-\mathrm{i}a(\sigma R)_{l}d_{l}^{k}\right)+\int_{\mathscr{U}}\biggl{(}f\big{(}x+g(u)\big{)}W_{1}(\sigma P_{1}u)^{\dagger}aW_{1}(\sigma P_{1}u)-af(x)\\ {}-\mathtt{1}_{\mathbb{S}}(u)\mathrm{i}f(x)\left[(P_{1}u)\mathtt{T}\sigma R,a\right]-\mathtt{1}_{\mathbb{S}}(u)a\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(u)\biggr{)}\nu(\mathrm{d}u);$ (41)
@@
However, the operators $L_{k}$ and $H(x)$ are unbounded, so that the construction of Secs. 2.2, 2.3 is purely formal. In any case (12), (13b) take the form
@@
The first equality implies that, under the physical probability $\mathbb{P}$, $\Pi(\mathrm{d}u,\mathrm{d}t)$ remains a Poisson measure with the same compensator. In the second equality the unbounded operators $R_{l}$ appear; so, to have $m_{k}(t)$ well defined depends on the choice of the coefficients $d_{l}^{k}$ and of the initial state. When $\left|m_{k}(t)\right|<+\infty$, from Eqs. (2.3.1), (19) we get the stochastic equation for the classical component
$\mathrm{d}X_{i}(t)=\left(\alpha^{0}_{i}\mathrm{d}t+\sum_{j=1}^{s}X_{j}(t_{-})Z^{00}_{ji}\mathrm{d}t+\sum_{k=1}^{d}a_{i}^{k}m_{k}(t)\right)\mathrm{d}t+\sum_{k=1}^{d}a_{i}^{k}\mathrm{d}\widehat{W}_{k}(t)\\ {}+\int_{\mathsf{U}_{0}}g_{i}(u)\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)+\int_{\mathscr{U}\setminus\mathsf{U}_{0}}g_{i}(u)\Pi(\mathrm{d}u,\mathrm{d}t),$ (44)
@@
$\mathrm{d}\langle\widehat{\sigma}_{t},a\rangle=\langle\widehat{\sigma}_{t_{-}},\mathcal{L}_{t}[a]\rangle\mathrm{d}t+\mathrm{i}\sum_{k=1}^{d}\sum_{l=1}^{2n}\langle\widehat{\sigma}_{t_{-}},R_{l}\overline{d^{k}_{l}}a-aR_{l}d^{k}_{l}\rangle\mathrm{d}\widehat{W}_{k}(t)-\sum_{k=1}^{d}m_{k}(t)\langle\widehat{\sigma}_{t_{-}},a\rangle\mathrm{d}\widehat{W}_{k}(t)\\ {}+\int_{\mathscr{U}}\langle\widehat{\sigma}_{t_{-}},W_{1}(P_{1}u)^{\dagger}aW_{1}(P_{1}u)-a\rangle\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t).$ (45)
In [13] it has been shown also that the adjoint of a quasi-free dynamical semigroup leaves invariant the sector of the state space based on classical densities with respect to Lebesgue measure, a fact which could not hold in the present case as discussed in Appendix A. So, the “quasi-free” requirement allows for unbounded quantum operators, but at the end it is a strong restriction on the possible structures.
@@ 4.1 A purely classical case
Let us take $\mathscr{H}=\mathbb{C}$ or, equivalently, take all the coefficients in Assumption 3 proportional to the identity $\mathds{1}$. Conditions (3) continue to hold for the functions $L_{k}(x)$ and $J(x,u)$. In this “classical” case we get from (7b), (14c)
@@
Then, there is no dynamics in the quantum component, as (19) gives $\mathrm{d}\widehat{\sigma}_{t}=0$. On the other side, the probability density $p_{t}$ has the non trivial evolution (14b) and it produces the new probability $\mathbb{P}$ (15), even without interaction with a quantum component.
Under $\mathbb{P}$, the classical process $X(t)$ satisfies the SDE (2.3.1), which looks like very similar to the original SDE (2) for a Markov process, apart from a peculiar difference: under $\mathbb{P}$ the jump noise $\Pi(\mathrm{d}u,\mathrm{d}t)$ is no more a Poisson measure, but a point process whose law depends on the solution $X(t)$ of the SDE itself, as discussed before (17). The advantage of working under the reference probability $\mathbb{Q}$ is to avoid the mathematical problems related to the involved SDE (2.3.1); the construction of the classical component $X(t)$ is split into the two standard SDEs (2) and (14b).
The effect of a non trivial probability density is apparent also from the formal generator (38), which becomes
$\mathcal{K}[a\otimes f](x)=a\Biggl{\{}\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\left(c_{i}(x)+2\sum_{k=1}^{d}a_{i}^{k}(x)\mathrm{\,Re\,}L_{k}(x)\right)+\sum_{i,j=1}^{s}\frac{1}{2}\,\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}a_{i}^{k}(x)a_{j}^{k}(x)\\ {}+\int_{\mathscr{U}}\left(\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\left|J(x,u)+1\right|^{2}-\mathtt{1}_{\mathsf{U}_{0}}(u)\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(x,u)\right)\nu(\mathrm{d}u)\Biggr{\}}.$
By comparing it with the formal generator (61), we see immediately the corrections due to the change of probability: $c_{i}(x)\to c_{i}(x)+2\sum_{k=1}^{d}a_{i}^{k}(x)\mathrm{\,Re\,}L_{k}(x)$ in the first term and $1\to\left|J(x,u)+1\right|^{2}$ in the jump term. While the first modification is only a redefinition of $c_{i}(x)$, the second modification is non trivial. Once again, the last modification is related to the fact that $\Pi(\mathrm{d}u,\mathrm{d}t)$ is no more a Poisson process, but a counting process with a random compensator $\left|J\big{(}X(t_{-}),u\big{)}+1\right|^{2}\nu(\mathrm{d}u)\mathrm{d}t$.
@@ 4.2 Jumps of discrete type
Then, Eqs.  (2), (12), (2.3) give
$\mathrm{d}X_{i}(t)=c_{i}\big{(}X(t_{-})\big{)}\mathrm{d}t+\sum_{k=1}^{d}a_{i}^{k}\big{(}X(t_{-})\big{)}\mathrm{d}W_{k}(t)+\sum_{k=1}^{l}g_{i}\big{(}X(t_{-}),k\big{)}\mathrm{d}N_{k}(t),\qquad i=1,\ldots,s,$ (46)
$\mathrm{d}\sigma_{t}={\mathcal{L}_{t}}_{*}[\sigma_{t_{-}}]\mathrm{d}t+\sum_{k=1}^{d}\left(L_{kt}\sigma_{t_{-}}+\sigma_{t_{-}}{L_{kt}}^{\dagger}\right)\mathrm{d}W_{k}(t)+\sum_{k=1}^{l}\left(J_{kt}\sigma_{t_{-}}J_{kt}^{\dagger}-\sigma_{t_{-}}\right)\mathrm{d}\widetilde{N}_{k}(t),$ (47)
$\mathcal{L}(x)_{*}[\rho]=-\mathrm{i}[H(x),\rho]+\sum_{k=1}^{d}\left(L_{k}(x)\rho L_{k}(x)^{\dagger}-\frac{1}{2}\left\{L_{k}(x)^{\dagger}L_{k}(x),\rho\right\}\right)\\ {}+\sum_{k=1}^{l}\lambda_{k}\left(J_{k}(x)\rho J_{k}(x)^{\dagger}-\frac{1}{2}\left\{J_{k}(x)^{\dagger}J_{k}(x),\rho\right\}\right),$ (48)
${\mathcal{L}_{t}}_{*}=\mathcal{L}\big{(}X(t_{-})\big{)}_{*},\qquad J_{kt}=J_{k}\big{(}X(t_{-})\big{)},\qquad\mathrm{d}\widetilde{N}_{k}(t)=\mathrm{d}N_{k}(t)-\lambda_{k}\mathrm{d}t.$
Under the reference probability $\mathbb{Q}$, $N_{k}(t)$, $k=1,\ldots,l$, are independent Poisson processes of intensities $\lambda_{k}$. Instead, under the physical probability $\mathbb{P}$, they become counting processes of intensities $\lambda_{k}I_{kt}$:
$I_{kt}=\langle\widehat{\sigma}_{t_{-}},\,J_{kt}^{\dagger}J_{kt}\rangle,\qquad\mathrm{d}\widehat{N}_{k}(t)=\mathrm{d}N_{k}(t)-\lambda_{k}I_{kt}.$
We assume to have $I_{kt}\neq 0$. Now, Eq. (19) gives
@@
We have written the linear SME (47) and the non-linear one (49) in strong form because only finite sums are involved and not integrals as in (12), (19).

### v2

@@ 3.3 A quasi-free dynamics
In [13] the structure of the most general quasi-free hybrid dynamical semigroup has been obtained. The “quasi-free” requirement means to map hybrid Weyl operators into a quantity proportional to hybrid Weyl operators [57]. In this case, the formal generator involves unbounded operators on the Hilbert space of the quantum component; so, the semigroup found in [13] is not in the class studied here. However, the comparison of the two formal generators gives hints to construct the stochastic formulation of the quasi-free dynamics.
@@
$\mathscr{U}=\Xi_{1}\oplus\Xi_{0},\qquad\Xi_{1}=\mathbb{R}^{2n},\qquad\Xi_{0}=\mathbb{R}^{s},\qquad\mathsf{U}_{0}=\mathbb{S}=\{u\in\mathscr{U}:\left|u\right|\leq 1\},$
$c_{i}(x)=\alpha^{0}_{i}+\sum_{j=1}^{s}x_{j}Z^{00}_{ji},\quad\alpha^{0}_{i}\in\mathbb{R}\quad Z^{00}_{ji}\in\mathbb{R},\quad i=1,\ldots,s,$
$b_{j}^{k}(x)=b_{j}^{k}\in\mathbb{R},\qquad g(x,u)=g(u)=P_{0}u\quad g_{i}(u)=u_{2n+i},$
@@
$\mathrm{d}X_{i}(t)=\alpha^{0}_{i}\mathrm{d}t+\sum_{j=1}^{s}X_{j}(t_{-})Z^{00}_{ji}\mathrm{d}t+\sum_{k=1}^{d}b_{i}^{k}\mathrm{d}W_{k}(t)\\ {}+\int_{\mathbb{S}}g_{i}(u)\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)+\int_{\mathscr{U}\setminus\mathbb{S}}g_{i}(u)\Pi(\mathrm{d}u,\mathrm{d}t),\qquad i=1,\ldots,s.$ (42)
@@
$H_{\rm q}=\beta^{\mathtt{T}}R+\frac{1}{2}\,R^{\mathtt{T}}P_{1}\sigma D^{11}\sigma^{\mathtt{T}}P_{1}R,\quad\beta\in\Xi_{1},$ (43d)
$H_{x}=x^{\mathtt{T}}P_{0}ZP_{1}R=\sum_{i=1}^{s}\sum_{j=1}^{2n}x_{i}Z^{01}_{ij}R_{j}.$ (43e)
@@
$\mathcal{K}[a\otimes f](x)=f(x)\Big{\{}\mathrm{i}[H_{\rm q}+H_{x},a]+\sum_{k=1}^{d}\sum_{r,l=1}^{2n}d^{k}_{l}\overline{d_{r}^{k}}\left((\sigma R)_{r}a(\sigma R)_{l}-\frac{1}{2}\left\{(\sigma R)_{r}(\sigma R)_{l},a\right\}\right)\\ {}+a\sum_{i=1}^{s}\left(\alpha^{0}_{i}+\sum_{j=1}^{s}x_{j}Z^{00}_{ji}\right)\frac{\partial f(x)}{\partial x_{i}}+a\sum_{i,j=1}^{s}\frac{1}{2}\,\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}b_{i}^{k}b_{j}^{k}\\ {}+\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\sum_{k=1}^{d}b_{i}^{k}\sum_{l=1}^{2n}\left(\mathrm{i}(\sigma R)_{l}\overline{d_{l}^{k}}a-\mathrm{i}a(\sigma R)_{l}d_{l}^{k}\right)+\int_{\mathscr{U}}\biggl{(}f\big{(}x+g(u)\big{)}W_{1}(\sigma P_{1}u)^{\dagger}aW_{1}(\sigma P_{1}u)-af(x)\\ {}-\mathtt{1}_{\mathbb{S}}(u)\mathrm{i}f(x)\left[(P_{1}u)\mathtt{T}\sigma R,a\right]-\mathtt{1}_{\mathbb{S}}(u)a\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(u)\biggr{)}\nu(\mathrm{d}u);$ (44)
@@
However, the operators $L_{k}$ and $H(x)$ are unbounded, so that the construction of Secs. 2.2, 2.3 becomes purely formal. In any case (10), (12) take the form
@@
The first equality implies that, under the physical probability $\mathbb{P}$, $\Pi(\mathrm{d}u,\mathrm{d}t)$ remains a Poisson measure with the same compensator. In the second equality the unbounded operators $R_{l}$ appear; so, to have $m_{k}(t)$ well defined depends on the choice of the coefficients $d_{l}^{k}$ and of the initial state. When $\left|m_{k}(t)\right|<+\infty$, from Eqs. (20), (21), (22) we get the stochastic equation for the classical component
$\mathrm{d}X_{i}(t)=\left(\alpha^{0}_{i}\mathrm{d}t+\sum_{j=1}^{s}X_{j}(t_{-})Z^{00}_{ji}\mathrm{d}t+\sum_{k=1}^{d}b_{i}^{k}m_{k}(t)\right)\mathrm{d}t+\sum_{k=1}^{d}b_{i}^{k}\mathrm{d}\widehat{W}_{k}(t)\\ {}+\int_{\mathsf{U}_{0}}g_{i}(u)\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)+\int_{\mathscr{U}\setminus\mathsf{U}_{0}}g_{i}(u)\Pi(\mathrm{d}u,\mathrm{d}t),$ (47)
@@
$\mathrm{d}\langle\widehat{\sigma}_{t},a\rangle=\langle\widehat{\sigma}_{t_{-}},\mathcal{L}_{t}[a]\rangle\mathrm{d}t+\mathrm{i}\sum_{k=1}^{d}\sum_{l=1}^{2n}\langle\widehat{\sigma}_{t_{-}},R_{l}\overline{d^{k}_{l}}a-aR_{l}d^{k}_{l}\rangle\mathrm{d}\widehat{W}_{k}(t)\\ {}-\sum_{k=1}^{d}m_{k}(t)\langle\widehat{\sigma}_{t_{-}},a\rangle\mathrm{d}\widehat{W}_{k}(t)+\int_{\mathscr{U}}\langle\widehat{\sigma}_{t_{-}},W_{1}(P_{1}u)^{\dagger}aW_{1}(P_{1}u)-a\rangle\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t).$ (48)
In [13] it has been shown also that the adjoint of a quasi-free dynamical semigroup leaves invariant the sector of the state space based on classical densities with respect to Lebesgue measure, a fact which could not hold in the general case as discussed in A. So, the “quasi-free” requirement allows for unbounded quantum operators, but at the end it is a strong restriction on the possible structures.
3.4 The master equation for hybrid states.
To have the explicit expression (41) of the formal generator can help in clarifying the various physical interactions involved in the hybrid dynamics and it allows to compare the trajectory approach with the approaches based on master equations. Let us exclude the cases discussed at the end of A and assume that the pre-adjoint ${\mathcal{T}_{t}}_{*}$ is well defined and that it leaves invariant $\mathscr{T}(\mathscr{H})\otimes L^{1}(\mathbb{R}^{s})$. Then, we can take as hybrid states the $x$-dependent trace class operators $\varrho(x)$ such that $\varrho(x)\geq 0$ and $\operatorname{Tr}\{\varrho(x)\}$ is a probability density with respect to the Lebesgue measure; then, we set $\varrho_{t}(x)={\mathcal{T}_{t}}_{*}[\varrho_{0}](x)$. To get the explicit master equation for $\varrho_{t}(x)$, we assume a smooth $x$-dependence and we simplify the jump contributions by taking $g(x,u)=g(u)$. By time-differentiation and using (41) and (12), we get the (heuristic) hybrid master equation
$\frac{\mathrm{d}\ }{\mathrm{d}t}\,\varrho_{t}(x)=\mathcal{K}_{*}[\varrho_{t}](x)=\mathcal{K}_{0}[\varrho_{t}](x)+\mathcal{K}_{\rm diff}[\varrho_{t}](x)+\mathcal{K}_{\rm jump}[\varrho_{t}](x),$ (49)
$\mathcal{K}_{0}[\varrho_{t}](x)=-\mathrm{i}[H(x),\varrho_{t}(x)]-\sum_{i=1}^{s}\frac{\partial\ }{\partial x_{i}}\bigl{(}c_{i}(x)\rho_{t}(x)\bigr{)},$ (50a)
$\mathcal{K}_{\rm diff}[\varrho_{t}](x)=\sum_{k=1}^{d}\Bigg{(}\frac{1}{2}\sum_{i,j=1}^{s}\frac{\partial^{2}\ \ }{\partial x_{i}\partial x_{j}}\left(b_{i}^{k}(x)b_{j}^{k}(x)\rho_{t}(x)\right)-\frac{1}{2}\left\{L_{k}(x)^{\dagger}L_{k}(x),\varrho_{t}(x)\right\}\\ {}+L_{k}(x)\varrho_{t}(x)L_{k}(x)^{\dagger}-\sum_{i=1}^{s}\frac{\partial\ }{\partial x_{i}}\Bigl{(}b_{i}^{k}(x)\bigl{(}L_{k}(x)\varrho_{t}(x)+\varrho_{t}(x)L_{k}(x)^{\dagger}\bigr{)}\Bigr{)}\Bigg{)},$ (50b)
$\mathcal{K}_{\rm jump}[\varrho_{t}](x)=\int_{\mathscr{U}}\nu(\mathrm{d}u)\Biggl{(}\Bigl{(}J\big{(}x-g(u),u\big{)}+\mathds{1}\Bigr{)}\varrho_{t}\big{(}x-g(u)\big{)}\Bigl{(}J\big{(}x-g(u),u\big{)}^{\dagger}+\mathds{1}\Bigr{)}\\ {}-\frac{1}{2}\left\{\Bigl{(}J(x,u)^{\dagger}+\mathds{1}\Bigr{)}\Bigl{(}J(x,u)+\mathds{1}\Bigr{)},\varrho_{t}(x)\right\}+\frac{1}{2}\left[J(x,u)^{\dagger}-J(x,u),\varrho_{(}x)\right]+\mathtt{1}_{\mathsf{U}_{0}}(u)\sum_{i=1}^{s}g_{i}(u)\,\frac{\partial\varrho_{t}(x)}{\partial x_{i}}\Biggr{)}.$ (50c)
This structure allows to compare, at a heuristic level, the present approach to other approaches based on master equations for hybrid states, so giving an idea of possible physical applications. References [4, 6, 31, 30] introduce master equations of jump type which are particular cases of (49) with $\mathcal{K}_{\rm diff}=0$. Then, the diffusive case is usually introduced by a suitable limit un the jump case; references [5, 4, 6, 31, 30, 3] give master equations of diffusive type which are particular cases of (49) with $\mathcal{K}_{\rm jump}=0$. Examples with two-level systems and harmonic oscillators are constructed. The most relevant application is the attempt of connecting gravity and quantum matter in the so called “Newtonian limit” [30, 31].
@@ 4.1 A purely classical case
Let us take $\mathscr{H}=\mathbb{C}$, which means that no quantum component is present. Now the quantities introduced in Assumption 3 become complex functions; conditions (4) continue to hold for the functions $L_{k}(x)$ and $J(x,u)$. In this “classical” case we get from (6), (15)
@@
Then, there is no dynamics in the quantum component, as (22) gives $\mathrm{d}\widehat{\sigma}_{t}=0$, while Eq. (10) reduces to (14). So, the probability density $p_{t}$ has the non trivial evolution (14) and it produces the new probability $\mathbb{P}$ (17), even without interaction with a quantum component.
Under $\mathbb{P}$, the classical process $X(t)$ satisfies the stochastic differential equation (20), which looks like very similar to the original stochastic equation (2) for a Markov process, apart from a peculiar difference: under $\mathbb{P}$ the jump noise $\Pi(\mathrm{d}u,\mathrm{d}t)$ is no more a Poisson measure, but a point process whose law depends on the solution $X(t)$ of the stochastic equation itself, as discussed before (19). The advantage of working under the reference probability $\mathbb{Q}$ is to avoid the mathematical problems related to the involved stochastic equation (20); the construction of the classical component $X(t)$ is split into the two standard stochastic differential equations (2) and (14).
The effect of a non trivial probability density is apparent also from the formal generator (41), which now is purely of classical type:
$\mathcal{K}[f](x)=\Biggl{\{}\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\left(c_{i}(x)+2\sum_{k=1}^{d}b_{i}^{k}(x)\mathrm{\,Re\,}L_{k}(x)\right)+\sum_{i,j=1}^{s}\frac{1}{2}\,\frac{\partial^{2}f(x)}{\partial x_{i}\partial x_{j}}\sum_{k=1}^{d}b_{i}^{k}(x)b_{j}^{k}(x)\\ {}+\int_{\mathscr{U}}\left(\left(f\big{(}x+g(x,u)\big{)}-f(x)\right)\left|J(x,u)+1\right|^{2}-\mathtt{1}_{\mathsf{U}_{0}}(u)\sum_{i=1}^{s}\frac{\partial f(x)}{\partial x_{i}}\,g_{i}(x,u)\right)\nu(\mathrm{d}u)\Biggr{\}}.$
By comparing it with the formal generator (68), we see immediately the corrections due to the change of probability: $c_{i}(x)\to c_{i}(x)+2\sum_{k=1}^{d}b_{i}^{k}(x)\mathrm{\,Re\,}L_{k}(x)$ in the first term and $1\to\left|J(x,u)+1\right|^{2}$ in the jump term. While the first modification is only a redefinition of $c_{i}(x)$, the second modification is non trivial. Once again, the last modification is related to the fact that $\Pi(\mathrm{d}u,\mathrm{d}t)$ is no more a Poisson process, but a counting process with a random compensator $\left|J\big{(}X(t_{-}),u\big{)}+1\right|^{2}\nu(\mathrm{d}u)\mathrm{d}t$.
To take a non trivial quantum component ($\mathscr{H}\neq\mathbb{C}$), but with all the coefficients in Assumption 3 proportional to the identity $\mathds{1}$, would give the same result on the classical component; indeed, this choice of the coefficients implies that the classical component does not receive any back-action from the quantum component. Also intermediate situations can arise with suitable choices of the coefficients. This example shows that a change of probability is not a purely quantum effect, but it may be due to the reaction of the quantum component on the classical one and to auto-interactions in the classical component.
@@ 4.2 Jumps of discrete type
In the classical case, described by (66), the integral over $\mathscr{U}$ has the role of giving rise to a continuity of types of jumps; moreover, the integral over $\mathsf{U}_{0}$ allows for ‘infinitely many small jumps’. These features are inherited by the hybrid case of Sec. 2. A simpler case, with sums instead of integrals, can help in the interpretation of the various terms.
@@
The pre-adjoint of the Liouville operator (12), acting on the trace class, becomes
$\mathcal{L}(x)_{*}[\rho]=-\mathrm{i}[H(x),\rho]+\sum_{k=1}^{d}\left(L_{k}(x)\rho L_{k}(x)^{\dagger}-\frac{1}{2}\left\{L_{k}(x)^{\dagger}L_{k}(x),\rho\right\}\right)\\ {}+\sum_{k=1}^{l}\lambda_{k}\left(J_{k}(x)\rho J_{k}(x)^{\dagger}-\frac{1}{2}\left\{J_{k}(x)^{\dagger}J_{k}(x),\rho\right\}\right).$ (51)
Moreover, as in (11), (15) we introduce $m_{k}(t)$ and we set
${\mathcal{L}_{t}}_{*}=\mathcal{L}\big{(}X(t_{-})\big{)}_{*},\qquad J_{kt}=J_{k}\big{(}X(t_{-})\big{)},\qquad I_{kt}=\langle\widehat{\sigma}_{t_{-}},\,J_{kt}^{\dagger}J_{kt}\rangle,$ (52)
Under the reference probability $\mathbb{Q}$, $N_{k}(t),\;k=1,\ldots,l,$ are independent Poisson processes of intensities $\lambda_{k}$; we set $\mathrm{d}\widetilde{N}_{k}(t)=\mathrm{d}N_{k}(t)-\lambda_{k}\mathrm{d}t$. Then, the hybrid dynamics is given by the coupled SDEs (2), (10), which now become
$\mathrm{d}X_{i}(t)=c_{i}\big{(}X(t_{-})\big{)}\mathrm{d}t+\sum_{k=1}^{d}b_{i}^{k}\big{(}X(t_{-})\big{)}\mathrm{d}W_{k}(t)+\sum_{k=1}^{l}g_{i}\big{(}X(t_{-}),k\big{)}\mathrm{d}N_{k}(t),$ (53)
$\mathrm{d}\sigma_{t}={\mathcal{L}_{t}}_{*}[\sigma_{t_{-}}]\mathrm{d}t+\sum_{k=1}^{d}\left(L_{kt}\sigma_{t_{-}}+\sigma_{t_{-}}{L_{kt}}^{\dagger}\right)\mathrm{d}W_{k}(t)+\sum_{k=1}^{l}\left(J_{kt}\sigma_{t_{-}}J_{kt}^{\dagger}-\sigma_{t_{-}}\right)\mathrm{d}\widetilde{N}_{k}(t).$ (54)
Under the physical probability $\mathbb{P}$, the $N_{k}(t)$ become counting processes of intensities $\lambda_{k}I_{kt}$ and we set $\mathrm{d}\widehat{N}_{k}(t)=\mathrm{d}N_{k}(t)-\lambda_{k}I_{kt}\mathrm{d}t$. We assume to have $I_{kt}\neq 0$. Now, the system of dynamical SDEs (20) and (22) becomes
$\mathrm{d}X_{i}(t)=\left(c_{i}\big{(}X(t_{-})\big{)}+\sum_{k=1}^{d}b_{i}^{k}\big{(}X(t_{-})\big{)}m_{k}(t)+\sum_{k=1}^{l}g_{i}\big{(}X(t_{-}),k\big{)}\lambda_{k}I_{kt}\right)\mathrm{d}t\\ {}+\sum_{k=1}^{d}b_{i}^{k}\big{(}X(t_{-})\big{)}\mathrm{d}\widehat{W}_{k}(t)+\sum_{k=1}^{l}g_{i}\big{(}X(t_{-}),k\big{)}\mathrm{d}\widehat{N}_{k}(t),$ (55)
@@
We have written the linear SME (54) and the non-linear one (56) in strong form because only finite sums are involved and not integrals as in (10), (22). Recall that the conditional mean of the driving increments, given the past frozen, vanishes; in symbols:
$\operatorname{\mathbb{E}}_{\mathbb{P}}\left[\mathrm{d}\widehat{W}_{k}(t)\Big{|}\mathscr{F}_{t_{-}}\right]=0,\qquad\operatorname{\mathbb{E}}_{\mathbb{P}}\left[\mathrm{d}\widehat{N}_{k}(t)\Big{|}\mathscr{F}_{t_{-}}\right]=0.$
This is because, under $\mathbb{P}$, $\widehat{W}_{k}(t)$ is a Wiener process and $\widehat{N}_{k}(t)$ is a compensated counting process; both processes are martingales.
Let us comment the dynamics (55) of the classical component. By the vanishing of the means of the increments of the driving processes, the terms in the second line of (55) can be interpreted as fluctuating forces; the first term is a continuous contribution and the second one gives also jumps. The quantum component contributes to these two terms only through the probability law of the processes $\widehat{N}_{k}(t)$. Instead, in the first line, a drift term appears. The first term in the drift does not depend on the quantum component and it represents some auto-interaction of the classical component; it could be generated by a classical Hamiltonian function. The two other terms in the drift can again depend on the classical component, as all the coefficients can contain $X(t_{-})$, but they depend also on the quantum state through the “quantum expectations” $m_{k}(t)$ and $I_{tk}$. They represent a back-reaction of the quantum component on the classical one. Let us stress that the presence of these back-reaction drifts needs the non-vanishing of the fluctuating forces, because they are connected by the coefficients $b^{k}(x)$ and $g(x,k)$: the classical evolution must be random if there is transfer of information from the quantum component. In the case without jumps, particular cases of (20), including the back-reaction, have been obtained also in other approaches, see for instance [3, (49)], [5, (7)]. Again similar non-linear back-reaction terms are obtained in completely different schemes in which the hybrid dynamics is obtained from microscopic models via Ehrenfest-like approximations [11].
A way to compare the trajectory approach to hybrid systems to approaches based on moments [12, 11] would be to compute the multi-time correlations of the classical component by using (55) and stochastic calculus, or by using the characteristic functions of the multi-time probabilities (33) and the characteristic operator given in (36). The computations of the multi-time moments in the case of continuous measurements and purely diffusive case are developed in [23, Sec. 4.3]. For the first moment it is enough to take the mean of (55); we get directly
$\frac{\mathrm{d}\ }{\mathrm{d}t}\operatorname{\mathbb{E}}_{\mathbb{P}}[X(t)]=\operatorname{\mathbb{E}}_{\mathbb{P}}\left[c\big{(}X(t)\big{)}\right]+\sum_{k=1}^{d}\operatorname{\mathbb{E}}_{\mathbb{P}}\left[b^{k}\big{(}X(t)\big{)}m_{k}(t)\right]+\sum_{k=1}^{l}\operatorname{\mathbb{E}}_{\mathbb{P}}\left[g\big{(}X(t),k\big{)}\lambda_{k}I_{kt}\right].$
Here we recognize immediately the terms discussed above: the mean of the auto-interaction of the classical component and of the two terms giving the back-action of the quantum component on the classical system.
About the dynamics (56) of the quantum component, we note the dependence on $X(t_{-})$ in the Liouville operator (51), (52). This gives the non-Markovian behaviour of the mean quantum dynamics, already commented after Eq. (25). When present, this $x$-dependence in (51) represents the action of the classical component on the quantum one, say a classical feedback or an external control [24, 28, 37, 58]. It can be shown that (56) preserves the pure states, as it is equivalent to the non-linear SSE recalled in Remark 1. Often, this equation is introduced as stochastic unravelling of the dissipative dynamics of the mean state [24, 3, 5].